OpenStudy (anonymous):

The London Eye (a ferris wheel) makes one rotation in approximately 30 minutes. What is the measure of the angle (in degrees) that a cart (spoke on the wheel) will rotate in 12 minutes?

OpenStudy (aaronq):

rate: \(\sf \dfrac{360^o}{30~min}\rightarrow 12^o/min\) degrees travelled in 12 min: \(\sf (12~min)(12^o/min)=144^o\)